40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 6840     Entry time: Wed Jun 20 18:09:23 2012
Author: yuta 
Type: Update 
Category: Locking 
Subject: both arms aligned, ITMX oplev centered 

[Jenne, Yuta]

We aligned FPMI. I also centered ITMX oplev because the light was not hitting on QPD.
Alignment procedure we took was;

1. Align Y arm to the Y end green(Y green trans to PSL is now 195 uW with Y end laser measured temperature 34.14 degC).
2. Aligned IR using PZT2 to Yarm(Now, TRY ~ 0.90).
3. Aligned ITMX monitoring AS spots.
4. Aligned X arm so that TRX maximize.
5. Fine adjusted both BS and X arm(Now, TRX ~ 0.82).

Beam spot position on ETMX looks a little too high & left (from ETMXF camera), but we will leave it until ASS scripts is fixed.

FPMIalignment2010620.png

ELOG V3.1.3-